OpenStudy (anonymous):

Let f(x) = 12 over the quantity of 4 x + 2 Find f(−1). −6 −2 2 6

Vocaloid (vocaloid):

x = -1 f(x) = 12/(4x + 2) = ?

OpenStudy (anonymous):

@Vocaloid 62x+1

OpenStudy (anonymous):

6/2x+1

Vocaloid (vocaloid):

uh, not quite... replace "x" with "-1" x = -1 f(x) = 12/(4x + 2) = 12/(4(-1)+2) = ?

OpenStudy (anonymous):

-6